OpenStudy (anonymous):

Choose the point-slope form of the equation below that represents the line that passes through the point (6, -3) and has a slope of 1/2. A. y - 6 = 1/2(x + 3) B. y = 1/2x - 6 C. y + 3 = 1/2(x - 6) D. x - 2y = 12

OpenStudy (anonymous):

This site is VERY helpful: https://www.desmos.com/calculator

OpenStudy (anonymous):

Do not give direct answers, and that is incorrect @shaadykhamlichi

OpenStudy (mrnood):

@101 +1

OpenStudy (shaadykhamlichi):

@iGreen someone needs help

OpenStudy (igreen):

Point-Slope Form: \(y - y_1 = m(x - x_1)\) Where \(x_1\) is the x-value of one of the points, \(y_1\) is the y-value of one of the points, and \(m\) is the slope.

OpenStudy (uri):

I can help you in this.

OpenStudy (uri):

the formula is y-y1=mx(x-x1)

OpenStudy (uri):

m=1/2

OpenStudy (igreen):

*y - y1 = m(x - x1)

OpenStudy (uri):



OpenStudy (igreen):

Here: \(x_1 = 6\) \(y_1 = -3\) \(m = \dfrac{1}{2}\)

OpenStudy (igreen):

Just plug it into the formula and you're done.